OpenStudy (anonymous):

Find the equation for the following: Passing the points A(3,0) and B(-2,-3). What are the steps to finding the answer?

OpenStudy (anonymous):

is it linear?

OpenStudy (anonymous):

It doesn't specify.

OpenStudy (anonymous):

first find the slope I guess by using this formula m=y2-y1/x2-x1

OpenStudy (anonymous):

okay so 3/5.

OpenStudy (anonymous):

use this m value and the coordinate of any of these points and substitute them in this equation : y=mx+b

OpenStudy (anonymous):

to get b

OpenStudy (anonymous):

then write the equation y=mx+b using the values you found for b and m

OpenStudy (anonymous):

would the final answer be y=3/5x-9/5

OpenStudy (anonymous):

yeah you can multiply both sides by 5 to get rid of the denominator

OpenStudy (anonymous):

and you can rearrange the equation to have all on one side like this : 3x-5y-9=0

OpenStudy (anonymous):

yeah that makes it look better, thank you for all your help :)

OpenStudy (anonymous):

you can check your answer by substituting one of the points in the equation just to make sure

OpenStudy (anonymous):

you are welcome